OpenStudy (anonymous):

Consider the following algorithm. for i is in {1, 2, 3, 4} do beep for j is in {1, 2, 3, 4} do beep for k is in {1, 2, 3} do for l is in {1, 2, 3, 4, 5} do beep for m is in {1, 2, 3, 4, 5, 6} do beep How many times does a beep statement get executed?

OpenStudy (anonymous):

\(4 + 4 \times 4 + 4 \times 4 \times 3 + 4 \times 4 \times 3 \times 5 + 4 \times 4 \times 3 \times 5 \times 6\)

OpenStudy (loser66):

@Alchemista what is it?

OpenStudy (anonymous):

I have to check my answer first.

OpenStudy (anonymous):

Yes I was right: \(4 + 4 \times 4 + 4 \times 4 \times 3 + 4 \times 4 \times 3 \times 5 + 4 \times 4 \times 3 \times 5 \times 6 =1748\) I tested it in python: x = 0 for _i in range(4): x += 1 for _i in range(4): x += 1 for _i in range(3): x += 1 for _i in range(5): x += 1 for _i in range(6): x += 1 print('Answer: {}'.format(x)) print('Prediction: {}'.format(4 + 4*4 + 4*4*3 + 4*4*3*5 + 4*4*3*5*6)) ------------------------------------------ db@vmbox ~ $ python test.py Answer: 1748 Prediction: 1748

OpenStudy (anonymous):

it is a wrong answer. but anyways thank you for your help.